OpenStudy (anonymous):

Find the line of symmetry for the parabola whose equation is y = 2x^2 - 4x + 1. Give the values of a, b, and c from the general form of the equation (2x + 1)(x - 2) = 0. What is the line of symmetry for the graph of y = -3x^2 + 12x - 11? What is the x-coordinate of the vertex of the parabola whose equation is y = 3x2 + 9x? Which of the following points lies on the graph of y = x^2 - 2x + 6? The line of symmetry of the parabola whose equation is y = ax2 - 4x + 3 is x = -2. What is the value of "a"?

OpenStudy (anonymous):

ind the line of symmetry for the parabola whose equation is y = 2x^2 - 4x + 1

OpenStudy (anonymous):

hey dad

OpenStudy (anonymous):

it is once again \(-\frac{b}{2a}\) with \(a=2,b=-4\)

OpenStudy (anonymous):

hello dear back from the rave already i see

OpenStudy (anonymous):

yeah i popped a molly can i have ice cream for breakfast?

OpenStudy (anonymous):

sure , but we are out of orange soda did you get \(x=1\) ?

OpenStudy (anonymous):

yeah, but would it be -? or - 1/2?

OpenStudy (anonymous):

oh whoops, nvm i see what i did wrong

OpenStudy (anonymous):

\[-\frac{b}{2a}=-\frac{-4}{2\times 2}=1\]

OpenStudy (anonymous):

i forgot the negatives cancelled

OpenStudy (anonymous):

right

OpenStudy (anonymous):

sorry dad still kinda rollin off the molly

OpenStudy (anonymous):

you will develop early onset Alzheimers like me

OpenStudy (anonymous):

Give the values of a, b, and c from the general form of the equation \[(2x + 1)(x - 2) = 0\]\[2 x^2-3 x-2=0\]

OpenStudy (anonymous):

\[2 x^2-3 x-2=0\\ ax^2+bx+c=0\]

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

yes dad

OpenStudy (anonymous):

What is the line of symmetry for the graph of y = -3x^2 + 12x - 11\[-\frac{b}{2a}=-\frac{12}{2\times (-3)}\]

OpenStudy (anonymous):

let me know when you get \(x=2\)

OpenStudy (anonymous):

i see -12/-6 = 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is the x-coordinate of the vertex of the parabola whose equation is \(y = 3x^2 + 9x\)? \[-\frac{b}{2a}=-\frac{9}{2\times 3}\]

OpenStudy (anonymous):

what does -9/6 simplify to?

OpenStudy (anonymous):

cancel a 3 top and bottom

OpenStudy (anonymous):

let me know when you get \(-\frac{3}{2}\)

OpenStudy (anonymous):

- 2/3?

OpenStudy (anonymous):

you got it upside down

OpenStudy (anonymous):

\[-\frac{9}{6}=-\frac{3}{2}\]

OpenStudy (anonymous):

it only gives me these 3 options, -3, -1 1/2, -2/3

OpenStudy (anonymous):

this is the equation \[y = 3x^2 + 9x\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ooh i see they want a mixed number

OpenStudy (anonymous):

\[-\frac{3}{2}=-1\tfrac{1}{2}\]

OpenStudy (anonymous):

ah i see

OpenStudy (anonymous):

Which of the following points lies on the graph of y = x^2 - 2x + 6 got choices?

OpenStudy (anonymous):

(-3,21), (-2,0), (-1,10)

OpenStudy (anonymous):

\[ (-3)^2 - 2(-3) + 6=21\]

OpenStudy (anonymous):

so first one is right didn't check the second one, hold on

OpenStudy (anonymous):

no second and third are wrong only the first one

OpenStudy (anonymous):

okay dad

OpenStudy (anonymous):

The line of symmetry of the parabola whose equation is\( y = ax^2 - 4x + 3\) is \(x = -2\). What is the value of "a"?

OpenStudy (anonymous):

\[-\frac{b}{2a}=-2\\ -\frac{-4}{2a}=-2\]

OpenStudy (anonymous):

or \[\frac{4}{2a}=-2\] so \[-4a=4\]

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

that it?

OpenStudy (anonymous):

gonna make a new one, thanks for all your help daddy!

OpenStudy (anonymous):

yw